# Stochastic surfing turbulent vorticity

### v1

@@ Methods
.2 Comparison between experimental and numerical results
[…]
@@
.3 Theoretical model
.3.1 Statistics of direction angle $\alpha$
When analyzing the mechanism of stochastic resonance, we employ a simplified model of the particle’s rotational dynamics. In this model, the governing equation is reformulated in terms of the phase lag angle $\beta$, under the assumption that the angle ($\alpha$) between the orientation direction ($\bm{n}$) and $z$-axis ($\bm{\hat{n}}$) satisfies $\alpha\to\pi/2$, which corresponds to $\kappa=\sin^{2}\alpha\to 1$. Based on this assumption, the equation simplifies to
@@
This simplification is justified by the statistical behavior of $\kappa=\sin^{2}{\alpha}$, whose probability density function is sharply peaked at 1 and rapidly decays for values smaller than 1. To validate this assumption, we plot the probability density distributions of $\sin^{2}\alpha$, denoted as $P(\sin^{2}\alpha)$, for different noise intensities, $\omega_{\eta}/\omega_{a}$, in both the subcritical ($\omega_{H}/\omega_{a}=0.1<1/2$) and supercritical ($\omega_{H}/\omega_{a}=1.1>1/2$) regimes. The results confirm that $\kappa$ is indeed strongly concentrated around 1 and the probability decreases rapidly for smaller values of $\alpha$ ($<\pi/2$), supporting the validity of our simplification.
.3.2 Linear stability analysis of the particle rotational dynamics in quiescent fluid
@@
.4 Waiting time statistics and equivalent noise intensity estimation
@@
The transition rate $\lambda$ varies with the noise intensity $\omega_{a}$, as shown in Supplementary Fig. 3d for different types of particles (light, neutral, and heavy). Specifically, $\lambda$ exhibits an exponential dependence on normalized potential barrier $\omega_{a}/\omega_{\eta}$, namely $\lambda\sim\exp(-\omega_{a}/\epsilon)$. This behavior is reminiscent of Kramers’ escape rate in a double-well potential subjected to thermal noise: given by $\lambda\sim e^{-\Delta U/\epsilon}$ [27, 25, 54], where $\Delta U$ is the potential barrier and $\epsilon$ represents the noise strength. In our case, the effective energy barrier between stable states of $\beta$ is proportional to $\omega_{a}$. Physically, $\epsilon$ quantifies the relative strength of turbulent fluctuations compared to the deterministic restoring forces acting on the particle. A larger $\epsilon$ implies a stronger stochastic influence, facilitating more frequent transitions between metastable states, while a lower $\epsilon$ implies that the system remains trapped in its metastable states for longer durations. This aligns with the earlier observation that particles of different densities explore distinct regions of the turbulence and thus different transition rates (slopes) are observed (Supplementary Fig. 3c). The estimated equivalent noise intensity shown in Supplementary Fig. 3e provides a direct means to quantify turbulence-induced noise intensity from observable transition dynamics. Within the investigated parameter space, turbulent fluctuations effectively act as a random forcing term that continuously perturbs the system, leading to memoryless escape events over the effective potential barriers. However, this picture may change if the particle enters the underdamped regime or has a finite size, where inertial effects should be taken into account [69, 74, 75].
@@
.5 Results of heavy and neutral particles
In addition to the simulation results presented in the main text on light particles in turbulence subjected to a rotating magnetic field, we also investigate the dynamics of particles with different densities. Specifically, we analyze neutral particles, characterized by a density contrast parameter $\beta_{p}=\frac{3}{1+2\rho_{p}/\rho_{f}}=1$ and heavy particles with $\beta_{p}=0.01$.

### v2

@@ Methods
I Experimental setup
I.1 Particle Characterization
In optical images as shown in Fig. 1, the particle surfaces may appear “rough.” This appearance originates from the contrast between the bright Styrofoam core and the darker magnetic coating, rather than actual geometric protrusions. Variations in paint deposition, the side of the particle in contact with the substrate during spraying, and small inconsistencies in the magnetic paint spraying angle result in non-uniform coating patterns. These patterns enhance optical contrast, creating the visual impression of surface roughness. However, the actual geometric irregularities are negligible compared to the particle diameter, and the surfaces remain effectively smooth on the hydrodynamic scale. Therefore, while such patterns may introduce minor perturbations to local flow, they do not significantly alter translational drag or rotational resistance beyond the effects already captured by particle size and magnetic torque.
The particles are designed to probe rotational dynamics under the combined influence of turbulence and a rotating magnetic field. Their size are comparable to the Kolmogorov scale, ensuring interactions with the smallest turbulent structures while remaining much smaller than the integral scale and Taylor microscale. Consequently, they cannot be regarded as point-like; however, finite-size effects on the rotation are limited, and the dominant particle–flow interactions occur at dissipative scales. The particles’ magnetic properties were characterized using a Superconducting Quantum Interference Device (SQUID) magnetometer (Quantum Design Inc.). Batch measurements yielded a mean volume susceptibility of $|\overline{\chi}_{V}|=0.136\pm 0.046$, while single-particle measurements revealed an average anisotropy of $\Delta\chi=0.011\pm 0.002$.
I.2 Magnetic field calibration
The rotating planar magnetic field was calibrated using a SENIS F3A magnetic-field-to-voltage transducer with an integrated 3D Hall probe. Across the frequency range of 1–50 Hz, the maximum magnetic flux density remained stable at $1.60\pm 0.02$ mT. The driving frequency is set to be 20 Hz for Fig. 2 and 10 Hz for Fig. 4, which illustrates the stochastic resonance. Further details of the field characterization are reported in our experimental study [61].
I.3 Flow field characterization
Our experiments are conducted in a Von Kármán-type turbulent flow, characterized by a Taylor-scale Reynolds number of up to $Re_{\lambda}=447$, ensuring fully developed homogeneous and isotropic turbulence in the measurement volume. It is noteworthy that the particle rotational dynamics is dominated by the turbulent dynamics at the small scales, i.e. the scales of vortex filaments. It is widely conjectured, motivated by numerical and experimental observations, that the small scales of turbulence are universal [86, 87, 88, 89, 90]. Therefore, it is expected that, provided the turbulence is fully developed, the strength of the large scale forcing, i.e., different values of the Reynolds number, poses negligible effects on the particle rotational dynamics. While the background flow is turbulent, the particle rotational dynamics are primarily governed by interactions with small-scale eddies. The particle rotational Reynolds number, $Re_{p}=\rho_{f}r^{2}|\omega_{p}-\omega_{\eta}|$, is estimated to be $\mathcal{O}(10^{-1})$ (below 1) based on typical maximum relative angular velocities, justifying the use of the laminar (Stokes) rotational drag coefficient, $\xi_{r}$ in out theoretical modelling. This ensures that viscous forces dominate inertial effects at the particle scale, providing an accurate leading-order description of particle-fluid interaction.
While the Von Kármán-type turbulent flow is not perfectly homogeneous and isotropic, a fully developed turbulent flow can be generated in the central region of the container with small-scale statistics close to local isotropy [47, 48, 62, 49, 63, 64, 50]. This region, measuring approximately $20$ mm $\times$ 2020 mm $\times$ 2020 mm, serves as the primary domain for measurements. Since our small particles ($r\sim 5\eta$) primarily interact with dissipative-scale velocity gradients in this central region, the observed stochastic resonance and symmetry breaking are governed by particle-scale turbulence and magnetic forcing rather than large-scale flow anisotropy.
The measured dissipation rate $\varepsilon$ is determined from the driving torque on the two impellers as $\varepsilon=T\Omega/(\rho V_{\text{w}})$ [91, 92], where $T$ is the sum of the driving torques from the impellers at the top and bottom of the water tank and $V_{\text{w}}=2.8\times 10^{-3}$ m3 is the volume of the water tank. The driving torques are measured using two strain gauge torque meters mounted on the motors. In our experiments, we investigated two turbulence intensities: (1) the weak-turbulence condition, for which the results are shown in Fig. 2(b–d) of the main text; and (2) the strong-turbulence condition, for which the results are presented in Supplementary Fig. 1(b). For the weak-turbulence case, the impellers rotate at 0.83 Hz. The corresponding mean energy dissipation rate is $\varepsilon=0.037\pm 0.002~\text{m}^{2}/\text{s}^{3}$, which yields a Taylor-scale Reynolds number of $Re_{\lambda}=398$, a Kolmogorov length scale of $\eta=0.072~\text{mm}$, and a Kolmogorov time scale of $\tau_{\eta}=5.20~\text{ms}$. For the strong-turbulence case, the impellers rotate at 1.21 Hz. The corresponding mean energy dissipation rate is $\varepsilon=0.075\pm 0.001~\text{m}^{2}/\text{s}^{3}$, which yields a Taylor-scale Reynolds number of $Re_{\lambda}=447$, a Kolmogorov length scale of $\eta=0.061~\text{mm}$, and a Kolmogorov time scale of $\tau_{\eta}=3.66~\text{ms}$.
II Comparison between experimental and numerical results
Previous studies have indeed explored the rotational dynamics of magnetically driven objects in quiescent environments [29, 30, 31, 14, 20, 21, 22, 24, 25], demonstrating phenomena such as phase locking and oscillatory motion. However, a fundamental distinction of our work lies in investigating these dynamics within the stochastic and chaotic environment of fully developed turbulence. This critical inclusion reveals a new class of phenomena, particularly the emergence of stochastic resonance where turbulent fluctuations play an active role in enhancing the particle response to an external periodic magnetic field. Furthermore, we uncover a novel symmetry-breaking mechanism for inducing net particle rotation in zero-mean turbulent vorticity.
In the main paper, we have presented selected comparisons of the PDFs of particle angular velocity, $\omega_{p,z}$, between experimental and numerical results (Fig. 2 in the main paper), for weak ($1.2~\text{mT}$), c intermediate ($1.4~\text{mT}$), and d strong ($1.6~\text{mT}$) magnetic intensities with constant rotational frequency of the magnetic field ($20~\text{Hz}$), to illustrate the key findings. The PDFs are obtained using 3000 (in experiments) and 512 (in simulations) particle trajectories. Here, we provide additional comparisons to further support our conclusions.
Supplementary Fig. 1 presents the probability density distributions of the particle angular velocity components $\omega_{p,i}$ ($i=x,y,z$) obtained from both experiments and numerical simulations. In Supplementary Fig. 1a (Supplementary Movie 9), where no magnetic field is applied, the particle rotation is purely driven by turbulence, exhibiting isotropic behavior. To facilitate a direct comparison with the experimental results (circles), the numerical results (squares) are flattened over all directions (i.e., reshape the angular velocity vector into a one-dimensional (1D) array while preserving the original data order). In Supplementary Fig. 1b (Supplementary Movie 10), when a magnetic field of $1.6~\text{mT}$ (same intensity as that of Fig. 2d in the main paper) is applied and the turbulence intensity is increased ($Re_{\lambda}=447$), the numerical and experimental results show good agreement, demonstrating that the numerical simulation is capable of capturing the essential physics of the experiment. Supplementary Fig. 1c further provides a phase diagram illustrating different particle rotational dynamic regimes, which is reproduced from Fig. 3m in the main paper. The experimental conditions corresponding to Supplementary Fig. 1b are marked as a green triangle. As shown in Supplementary Figures 1a and b, the simulation results agree well with the experimental results. The slight discrepancies are due to the polydispersity of the particles. Specifically, these particles exhibit anisotropy in three key aspects: (1) Shape: The particles are nearly spherical, with an average diameter of $0.762\pm 0.066$ mm. However, minor deviations from perfect sphericity exist due to natural variations in the manufacturing process. These variations can introduce small differences in rotational dynamics, particularly when particles experience complex hydrodynamic and magnetic interactions in turbulence. (2) Density: The mean density of the magnetic particles is measured to be $\rho_{\text{particle}}=0.208\pm 0.014$ kg/m3. This value is significantly lower than that of water, indicating that the particles remain buoyant in the experimental fluid. The density variation across different particles arises from heterogeneities in the composition of the Styrofoam core and the magnetic coating. (3) Magnetic Properties: The magnetic characteristics of the particles are inherently non-uniform due to the way the magnetic paint adheres to the surface during fabrication. Unlike ideal dipoles, these particles exhibit variability in three crucial aspects: the magnitude of the magnetic moment $|\bm{m}|$, the orientation of the magnetic moment relative to the particle’s geometric axes, and the spatial distribution of magnetization across the surface. This non-uniformity means that different particles respond to an externally applied magnetic field in distinct ways, leading to variations in their rotational dynamics. Furthermore, the magnetization patterns on the particle surface serve as markers for tracking their rotational motion in experiments. All these three aspects of particle polydispersity can introduce minor differences between the simulation results w.r.t the experimental observations. However, these differences remain small, and overall, the simulations show good agreement with the experiments.
@@ II Comparison between experimental and numerical results
III Theoretical model
III.1 Statistics of direction angle
When analyzing the mechanism of stochastic resonance, we employ a simplified model of the particle rotational dynamics. In this model, the governing equation is reformulated in terms of the phase lag angle $\beta$, under the assumption that the angle ($\alpha$) between the orientation direction ($\bm{n}$) and $z$-axis ($\bm{\hat{n}}$) satisfies $\alpha\to\pi/2$, which corresponds to $\kappa=\sin^{2}\alpha\to 1$. Based on this assumption, the equation simplifies to
@@ III.1 Statistics of direction angle
This simplification is justified by the statistical behavior of $\kappa=\sin^{2}{\alpha}$, whose probability density function is sharply peaked at 1 and rapidly decays for values smaller than 1. To validate this assumption, we plot the probability density distributions of $\sin^{2}\alpha$, denoted as $P(\sin^{2}\alpha)$, for different noise intensities, $\omega_{\eta}/\omega_{a}$, in both the subcritical ($\omega_{H}/\omega_{a}=0.1<1/2$) and supercritical ($\omega_{H}/\omega_{a}=1.1>1/2$) regimes. The results confirm that $\kappa$ is indeed strongly concentrated around 1 and the probability decreases rapidly for smaller values of $\alpha$ ($\pi/2$), supporting the validity of our simplification.
III.2 Overdamped approximation justification
To justify the validity of the overdamped model, we carefully assessed inertial effects. For a spherical particle, the rotational response time can be estimated as $\tau_{\text{rot}}=\frac{I}{\xi_{r}}$, where $I=\frac{2}{5}\mathcal{V}_{p}\rho_{p}r^{2}$ is the particle moment of inertia, and $\xi_{r}=8\pi\mu r^{3}$ is the rotational drag coefficient. Substituting these into the equation for $\tau_{\text{rot}}$ simplifies to $\tau_{\text{rot}}=\frac{\rho_{p}r^{2}}{15\mu}$. Using our experimental parameters of $r_{p}=0.381\text{ mm}$, $\rho_{p}=208\text{ kg/m}^{3}$, $\mu\approx 1.0\times 10^{-3}\text{ Pa}\cdot\text{s}$, and $\tau_{\eta}=7.56\text{ ms}$. The rotational response time is $\tau_{\text{rot}}=2.01\text{ ms}$. We observe that $\tau_{\text{rot}}$ is approximately $2.01/7.56=0.266$ or about one-quarter of $\tau_{\eta}$, indicating that the particle rotational dynamics are significantly faster than the characteristic changes in the flow vorticity. Therefore, the particle can respond quickly to the flow vorticity, making the assumption of an overdamped regime a reasonable first approximation. Note that our experimental frequencies are chosen to ensure the particle can follow the driving field within the overdamped regime, i.e., the frequency of the magnetic field $\omega_{H}$ is slower than $1/\tau_{rot}\approx 2000~\text{Hz}$, so that the particle rotational motion will remain in the overdamped regime.
III.3 Neglecting rotation-translation coupling
In the model of particle angular dynamics, we neglect rotation–translation coupling by assuming that the particle center of mass and hydrodynamic center are approximately aligned. This assumption is justified by the nearly spherical shape, uniform mass distribution, and small size of the particles (compared to the Kolmogorov scale), which collectively minimize translation-induced hydrodynamic torque. Consequently, the hydrodynamic torque induced by translational motion is expected to be negligible. Even if weak coupling exists, its effect can be interpreted as an additional source of rotational noise that is uncorrelated with the external magnetic field. This may slightly modify the amplitude of angular fluctuations or slightly shift the resonance frequency, but it does not alter the underlying mechanism of stochastic resonance. Since the emergence of stochastic resonance in our system fundamentally arises from the balance between external magnetic forcing and intrinsic rotational noise, the presence of small translation–rotation coupling does not qualitatively affect our main conclusions.
III.4 Validity of the point-particle approximation
Despite neglecting finite-size hydrodynamic effects, the point-particle approximation offers a leading-order description of particle–fluid interaction and is widely adopted in studies of particle-laden flows [78, 81]. In our system, the particle radius is comparable to the Kolmogorov length scale, ensuring locally linear flow gradients that primarily govern rotational dynamics via local vorticity. This, coupled with a dominant external magnetic torque and a dilute suspension (volume fraction of $\Phi_{V}=0.17\%$ that minimizes particle-particle interactions [93]), provides a robust physical justification for our model. Crucially, the good agreement between our simulations and experimental measurements further validates the point-particle approximation for this system.
III.5 Linear stability analysis of the particle rotational dynamics in quiescent fluid
@@
IV Waiting time statistics and equivalent noise intensity estimation
@@
The transition rate $\lambda$ varies with the noise intensity $\omega_{a}$, as shown in Supplementary Fig. 3d for different types of particles (light, neutral, and heavy). Specifically, $\lambda$ exhibits an exponential dependence on the normalized potential barrier $\omega_{a}/\omega_{\eta}$, namely $\lambda\sim\exp(-\omega_{a}/\epsilon)$. This behavior is reminiscent of Kramers’ escape rate in a double-well potential subjected to thermal noise: given by $\lambda\sim e^{-\Delta U/\epsilon}$ [36, 34, 66], where $\Delta U$ is the potential barrier and $\epsilon$ represents the noise strength. In our case, the effective energy barrier between stable states of $\beta$ is proportional to $\omega_{a}$. Physically, $\epsilon$ quantifies the relative strength of turbulent fluctuations compared to the deterministic restoring forces acting on the particle. A larger $\epsilon$ implies a stronger stochastic influence, facilitating more frequent transitions between metastable states, while a lower $\epsilon$ implies that the system remains trapped in its metastable states for longer durations. This aligns with the earlier observation that particles of different densities explore distinct regions of the turbulence and thus different transition rates (slopes) are observed (Supplementary Fig. 3c). The estimated equivalent noise intensity shown in Supplementary Fig. 3e provides a direct means to quantify turbulence-induced noise intensity from observable transition dynamics. Within the investigated parameter space, turbulent fluctuations effectively act as a random forcing term that continuously perturbs the system, leading to memoryless escape events over the effective potential barriers. However, this picture may change if the particle enters the underdamped regime or has a finite size, where inertial effects should be taken into account [81, 94, 95].
@@
V Results of heavy and neutral particles
In addition to the simulation results presented in the main paper on light particles in turbulence subjected to a rotating magnetic field, we also investigate the dynamics of particles with different densities. Specifically, we analyze neutral particles, characterized by a density contrast parameter $\beta_{p}=\frac{3}{1+2\rho_{p}/\rho_{f}}=1$ and heavy particles with $\beta_{p}=0.01$.